# Dichotomous Diffusion Policy Optimization

### v1

@@ 4.1 Experiments on RL Benchmarks
Experimental setup. We evaluate our approach on two commonly-used benchmarks, OGBench (ogbench_park2025) task suite and ExORL (yarats2022exorl) benchmark. OGBench provides challenging robotic locomotion and manipulation tasks, including complex whole-body humanoid control, maze navigation, and object manipulation. We use the default dataset collected by RND (burda2018explorationrandomnetworkdistillation), including tasks in complex high-dimensional state-based domains: Walker, Quadruped, Jaco, and Cheetah. Our evaluation encompasses 30 tasks across 6 domains on OGBench and 9 tasks across 4 domains on ExORL for offline learning, totaling 39 tasks. Finally, we select 4 default tasks across 4 domains on OGBench for offline-to-online validation. Further details are provided in Appendix D.1.
@@
  • Gaussian policy. Standard RL uses Gaussian policies by default. In comparison with standard methods, we select 1) IQL (kostrikov2021offline): a typical weighted regression offline RL method. 2) ReBRAC (tarasov2023revisiting): an effective behavior-regularized actor-critic approach incorporates several specific designs tailored for offline learning.
  • Diffusion/Flow policy. We also include offline RL baselines built on diffusion or flow policies according to the following learning strategies: 1) IDQL (hansen2023idql) and IFQL (fql_park2025): both approaches employ expectile regression for value learning and utilize imitation pre-trained diffusion or flow models with rejection sampling during inference. 2) FQL (fql_park2025): a behavior-regularized actor-critic variant that uses flow policy distillation and shows strong performance on OGBench. 3) CFGRL (frans2025diffusion): a recently proposed policy improvement framework relies on classifier-free guidance, which uses high-quality actions for conditional policy training and unconditional behavior cloning.
For the offline RL setting, we compare our approach with IQL, ReBRAC, CFGRL, IFQL, and FQL on the ExORL benchmark. We also include a variant, DIPOLE w/o rs, which does not use rejection sampling during inference for clear comparison. For IFQL, we utilizes the default hyperparameters, and for FQL, we select the hyperparameter $\alpha$ reported in previous work (fql_park2025) with optimal performance in ExORL. Additionally, we compare with IQL, ReBRAC, IDQL, IFQL, and FQL on OGBench to demonstrate our method’s effectiveness against state-of-the-art approaches in challenging benchmark.
@@
Evaluation results. We evaluate the performance of each method after a fixed number of offline updates for the offline RL setting. Specifically, we report the average return on ExORL and the success rate on OGBench, following the standard evaluation assessment methods  (yarats2022exorl; ogbench_park2025). For the offline-to-online evaluation, we assess the final performance on a fixed number of online updates following the formal offline pretraining stage. All evaluations are averaging over 8 random seeds, with $\pm$ indicating standard deviations in tables. We present our results by answering the following questions:
@@ 4.2 Experiments on Autonomous Driving Benchmark
Our method is evaluated on the large-scale real-world autonomous driving benchmark NAVSIM (Dauner2024NEURIPS) using closed-loop assessment. Following the official evaluation protocol, we report the PDM score (higher indicates better performance), which aggregates five key metrics: NC (no-collision rate), DAC (drivable area compliance), TTC (time-to-collision safety), Comfort (acceleration/jerk constraints), and EP (ego progress). All methods are tested under the official closed‑loop simulator, and results are averaged over the public test split. We also consider an RL application scenario where RL can be applied in human take-over situations or complex environments lacking ground-truth supervision. To address this, we provide a variant of our model trained on the test split without using any ground-truth.
@@
We select several baselines: 1) UniAD (hu2023planning): integrates multiple auxiliary tasks such as tracking, mapping, prediction, and occupancy prediction using transformer blocks, and employs latent representations for planning. 2) PARA-Drive (weng2024drive): adopts a parallel architecture design compared to UniAD. 3) Transfuser (chitta2023transfuser): fuses image and LiDAR information through a dual-branch architecture and incorporates detection and BEV semantic maps for auxiliary supervision. Its latent variant, LFT, replaces LiDAR inputs with learnable embeddings. 4) Hydra-MDP (li2024hydra): winner of the CVPR2024 Challenge, which uses trajectory anchors and a learned reward model for anchor selection. Moreover, we also consider baselines where the agent either maintains its current state or uses a simple MLP for trajectory regression. For our imitation pre-trained VLA model, which directly generates trajectories without post-processing, we denote it as DP-VLA. When fine-tuned with DPPO(ren2024diffusion), we refer to them as DP-VLA w/ DPPO. When fine-tuned with our RL algorithm, we refer to it as DP-VLA w/ DIPOLE. As mentioned above, we also provide two variants trained on the navtrain and navtest splits.
@@ 5 Related Work
Reinforcement fine-tuning of diffusion models remains challenging due to their multi-step diffusion process, primarily in terms of learning stability and computational efficiency. A brute-force solution involves directly optimizing the reward via gradient backpropagation. ReFL (xu2023imagereward) optimizes human preference scores for image generation by backpropagating gradients at specific single steps during the reverse process. DRaFT (clark2023directly) extends this approach by applying gradient optimization across multiple steps at the end of the reverse process. Such methods are widely used in motion generation (karunratanakul2024optimizing), image generation (prabhudesai2023aligning), and decision-making tasks (wangdiffusion), but suffer from instability due to noisy gradient backpropagation during the denoising process. Some methods avoid gradient computation and instead search for the optimal noise to maximize reward, a strategy referred to as inference-time scaling (hansen2023idql; ma2025inference; singhal2025general). Recent approaches also utilize RL to directly search for the best noise (wagenmaker2025steering). In both cases, performance remains constrained by the capabilities of the pre-trained model. Moreover, DDPO (black2023training) treats each noise step as a Gaussian distribution, enabling likelihood estimation and optimization via the REINFORCE (mohamed2020monte) algorithm. DPPO (ren2024diffusion) optimizes this approach and extends it to multi-step MDPs using PPO (schulman2017proximal) for policy improvement. These methods rely on Gaussian approximations that require sufficiently small sampling steps, resulting in inefficient training. Some methods (lee2023aligning; kang2023efficient; zheng2024safe; maefficient; zheng2025towards) use KL-regularized RL (kostrikov2021offline; peng2019advantage), whose solution leads to a simple weighted regression loss. However, these approaches often face a trade-off between greediness and stability.
@@ Appendix B Theoretical Interpretations
We define our problem under the reinforcement learning problem presented as a Markov Decision Process (MDP) (sutton1998reinforcement) given by $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$, which comprises a state space $\mathcal{S}$, an action space $\mathcal{A}$, a state transition $\mathcal{S}$, a reward function $r$ and a discount factor $\gamma$. In this setting, a policy is a probability distribution of actions conditioned on a state. In addition, we assume that all policies induce an irreducible Markov Chain, with any two states reachable from each other by a sequence of transitions that have positive probability. Our goal is to find a policy $\pi$ that maximizes a predefined action evaluation criteria $G$, constrained on a reference policy.
@@ D.1 Experimental details
In this section, we provide the experimental details, including benchmarks, datasets, and tasks. Our experiments span two primary benchmarks: ExORL (yarats2022exorl) and OGBench (ogbench_park2025)
ExORL. ExORL consists of datasets collected by multiple unsupervised RL agents (laskin2021urlbunsupervisedreinforcementlearning) on the DeepMind Control Suite (tassa2018deepmindcontrolsuite). We utilize datasets collected by unsupervised RL algorithms RND (burda2018explorationrandomnetworkdistillation) across four domains (Walker, Jaco, Quadruped, and Cheetah). For each environment, we use the full dataset with all transitions from each dataset.
@@ D.2 Implementation details
We implement DIPOLE in JAX on top of FQL (fql_park2025) and CFGRL (frans2025diffusion).
@@
Evaluation. We report the average return on ExORL and the success rate on OGBench, following the standard evaluation assessment methods  (yarats2022exorl; ogbench_park2025). For offline RL performance evaluation, we fix the gradient to be 1M and report the final score. For the offline-to-online RL performance evaluation, we report both the 1M offline score and the 1M online score.
@@ E.1 Model Architecture
We employ the pretrained Florence-2-large model (xiao2024florence) as the visual-language encoder, paired with a 475M-parameter Diffusion Transformer as the action decoder. The visual input comprises images from Front, Front-Left, and Front-Right perspectives, while the language input consists of driving commands provided by the dataset. Encoder output tokens are processed by the action decoder through a cross-attention block, which ultimately generates the predicted trajectory.
@@ E.2 Training Procedure
The training process consists of two phases: pretraining and reinforcement learning fine-tuning. In the pretraining phase, we utilize trainval frames from the NAVSIM dataset to jointly train the encoder and decoder using a diffusion loss objective. During the RL fine-tuning phase, the encoder is frozen, and two Low-Rank Adaptation (LoRA) adapters—a positive adapter and a negative adapter—are incorporated into every linear projection of the attention and MLPs. The NavSim benchmark’s PDMS score serves as the direct optimization target. Every 10 epochs, the replay buffer is cleared, and new model rollout trajectories and corresponding rewards are collected based on one epoch of data samples. For each data sample, the model generates $g$ trajectories, which are used to train the LoRA adapters over the subsequent 9 epochs. For each trajectory, we obtain its PDMS score vector $\textbf{r}=\{r_{1},r_{2},\dots,r_{g}\}$, enabling estimation of the advantage function (shao2024deepseekmath):
@@ E.3 Inference and Evaluation
During inference, the DP-VLA encoder extracts features from the input images and driving commands. The denoising process is solved using the DPM-Solver(lu2022dpm), with the action decoder iteratively predicting the denoised trajectory over 10 steps to produce the final clean trajectory. For evaluation on the NAVSIM benchmark, the proposed trajectory is fed into an LQR tracker and dynamics model to compute the posterior trajectory. The final PDM score is derived from this posterior trajectory, satisfying the benchmark’s evaluation criteria (Dauner2024NEURIPS).

### v2

@@ 4.1 Experiments on RL Benchmarks
Experimental setup. We evaluate our approach on two commonly-used benchmarks, OGBench (Park et al., 2025a) task suite and ExORL (Yarats et al., 2022) benchmark. OGBench provides challenging robotic locomotion and manipulation tasks, including complex whole-body humanoid control, maze navigation, and object manipulation. We use the default dataset collected by RND (Burda et al., 2019), including tasks in complex high-dimensional state-based domains: Walker, Quadruped, Jaco, and Cheetah. Our evaluation encompasses 30 tasks across 6 domains on OGBench and 9 tasks across 4 domains on ExORL for offline learning, totaling 39 tasks. Finally, we select 4 default tasks across 4 domains on OGBench for offline-to-online validation. Further details are provided in Appendix D.1.
@@
  • Gaussian policy. Standard RL uses Gaussian policies by default. In comparison with standard methods, we select 1) IQL (Kostrikov et al., 2022): a typical weighted regression offline RL method. 2) ReBRAC (Tarasov et al., 2023): an effective behavior-regularized actor-critic approach incorporates several specific designs tailored for offline learning.
  • Diffusion/Flow policy. We also include offline RL baselines built on diffusion or flow policies according to the following learning strategies: 1) IDQL (Hansen-Estruch et al., 2023) and IFQL (Park et al., 2025b): both approaches employ expectile regression for value learning and utilize imitation pre-trained diffusion or flow models with rejection sampling during inference. 2) FQL (Park et al., 2025b): a behavior-regularized actor-critic variant that uses flow policy distillation and shows strong performance on OGBench. 3) CFGRL (Frans et al., 2025): a recently proposed policy improvement framework relies on classifier-free guidance, which uses high-quality actions for conditional policy training and unconditional behavior cloning.
For the offline RL setting, we compare our approach with IQL, ReBRAC, CFGRL, IFQL, and FQL on the ExORL benchmark. We also include a variant, DIPOLE w/o rs, which does not use rejection sampling during inference for clear comparison. For IFQL, we utilizes the default hyperparameters, and for FQL, we select the hyperparameter $\alpha$ reported in previous work (Park et al., 2025b) with optimal performance in ExORL. Additionally, we compare with IQL, ReBRAC, IDQL, IFQL, and FQL on OGBench to demonstrate our method’s effectiveness against state-of-the-art approaches in challenging benchmark.
@@
Evaluation results. We evaluate the performance of each method after a fixed number of offline updates for the offline RL setting. Specifically, we report the average return on ExORL and the success rate on OGBench, following the standard evaluation assessment methods  (Yarats et al., 2022; Park et al., 2025a). For the offline-to-online evaluation, we assess the final performance on a fixed number of online updates following the formal offline pretraining stage. All evaluations are averaging over 8 random seeds, with $\pm$ indicating standard deviations in tables. We present our results by answering the following questions:
@@ 4.2 Experiments on Autonomous Driving Benchmark
Our method is evaluated on the large-scale real-world autonomous driving benchmark NAVSIM (Dauner et al., 2024) using closed-loop assessment. Following the official evaluation protocol, we report the PDM score (higher indicates better performance), which aggregates five key metrics: NC (no-collision rate), DAC (drivable area compliance), TTC (time-to-collision safety), Comfort (acceleration/jerk constraints), and EP (ego progress). All methods are tested under the official closed‑loop simulator, and results are averaged over the public test split. We also consider an RL application scenario where RL can be applied in human take-over situations or complex environments lacking ground-truth supervision. To address this, we provide a variant of our model trained on the test split without using any ground-truth.
@@
We select several baselines: 1) UniAD (Hu et al., 2023): integrates multiple auxiliary tasks such as tracking, mapping, prediction, and occupancy prediction using transformer blocks, and employs latent representations for planning. 2) PARA-Drive (Weng et al., 2024): adopts a parallel architecture design compared to UniAD. 3) Transfuser (Chitta et al., 2023): fuses image and LiDAR information through a dual-branch architecture and incorporates detection and BEV semantic maps for auxiliary supervision. Its latent variant, LFT, replaces LiDAR inputs with learnable embeddings. 4) Hydra-MDP (Li et al., 2024): winner of the CVPR2024 Challenge, which uses trajectory anchors and a learned reward model for anchor selection. Moreover, we also consider baselines where the agent either maintains its current state or uses a simple MLP for trajectory regression. For our imitation pre-trained VLA model, which directly generates trajectories without post-processing, we denote it as DP-VLA. When fine-tuned with DPPO(Ren et al., 2025), we refer to them as DP-VLA w/ DPPO. When fine-tuned with our RL algorithm, we refer to it as DP-VLA w/ DIPOLE. As mentioned above, we also provide two variants trained on the navtrain and navtest splits.
@@ 5 Related Work
Reinforcement fine-tuning of diffusion models remains challenging due to their multi-step diffusion process, primarily in terms of learning stability and computational efficiency. A brute-force solution involves directly optimizing the reward via gradient backpropagation. ReFL (Xu et al., 2023b) optimizes human preference scores for image generation by backpropagating gradients at specific single steps during the reverse process. DRaFT (Clark et al., 2023) extends this approach by applying gradient optimization across multiple steps at the end of the reverse process. Such methods are widely used in motion generation (Karunratanakul et al., 2024), image generation (Prabhudesai et al., 2023), and decision-making tasks (Wang et al., 2022), but suffer from instability due to noisy gradient backpropagation during the denoising process. Some methods avoid gradient computation and instead search for the optimal noise to maximize reward, a strategy referred to as inference-time scaling (Hansen-Estruch et al., 2023; Ma et al., 2025b; Singhal et al., 2025). Recent approaches also utilize RL to directly search for the best noise (Wagenmaker et al., 2025). In both cases, performance remains constrained by the capabilities of the pre-trained model. Moreover, DDPO (Black et al., 2024b) treats each noise step as a Gaussian distribution, enabling likelihood estimation and optimization via the REINFORCE (Mohamed et al., 2020) algorithm. DPPO (Ren et al., 2025) optimizes this approach and extends it to multi-step MDPs using PPO (Schulman et al., 2017) for policy improvement. These methods rely on Gaussian approximations that require sufficiently small sampling steps, resulting in inefficient training. Some methods (Lee et al., 2023; Kang et al., 2023; Zheng et al., 2024; Ma et al., 2025a; Zheng et al., 2025b) use KL-regularized RL (Kostrikov et al., 2022; Peng et al., 2019), whose solution leads to a simple weighted regression loss. However, these approaches often face a trade-off between greediness and stability.
@@ Appendix B Theoretical Interpretations
We define our problem under the reinforcement learning problem presented as a Markov Decision Process (MDP) (Sutton et al., 1998) given by $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$, which comprises a state space $\mathcal{S}$, an action space $\mathcal{A}$, a state transition $\mathcal{S}$, a reward function $r$ and a discount factor $\gamma$. In this setting, a policy is a probability distribution of actions conditioned on a state. In addition, we assume that all policies induce an irreducible Markov Chain, with any two states reachable from each other by a sequence of transitions that have positive probability. Our goal is to find a policy $\pi$ that maximizes a predefined action evaluation criteria $G$, constrained on a reference policy.
@@ D.1 Experimental details
In this section, we provide the experimental details, including benchmarks, datasets, and tasks. Our experiments span two primary benchmarks: ExORL (Yarats et al., 2022) and OGBench (Park et al., 2025a)
ExORL. ExORL consists of datasets collected by multiple unsupervised RL agents (Laskin et al., 2021) on the DeepMind Control Suite (Tassa et al., 2018). We utilize datasets collected by unsupervised RL algorithms RND (Burda et al., 2019) across four domains (Walker, Jaco, Quadruped, and Cheetah). For each environment, we use the full dataset with all transitions from each dataset.
@@ D.2 Implementation details
We implement DIPOLE in JAX on top of FQL (Park et al., 2025b) and CFGRL (Frans et al., 2025).
@@
Evaluation. We report the average return on ExORL and the success rate on OGBench, following the standard evaluation assessment methods  (Yarats et al., 2022; Park et al., 2025a). For offline RL performance evaluation, we fix the gradient to be 1M and report the final score. For the offline-to-online RL performance evaluation, we report both the 1M offline score and the 1M online score.
@@ E.1 Model Architecture
We employ the pretrained Florence-2-large model (Xiao et al., 2024) as the visual-language encoder, paired with a 475M-parameter Diffusion Transformer as the action decoder. The visual input comprises images from Front, Front-Left, and Front-Right perspectives, while the language input consists of driving commands provided by the dataset. Encoder output tokens are processed by the action decoder through a cross-attention block, which ultimately generates the predicted trajectory.
@@ E.2 Training Procedure
The training process consists of two phases: pretraining and reinforcement learning fine-tuning. In the pretraining phase, we utilize trainval frames from the NAVSIM dataset to jointly train the encoder and decoder using a diffusion loss objective. During the RL fine-tuning phase, the encoder is frozen, and two Low-Rank Adaptation (LoRA) adapters—a positive adapter and a negative adapter—are incorporated into every linear projection of the attention and MLPs. The NavSim benchmark’s PDMS score serves as the direct optimization target. Every 10 epochs, the replay buffer is cleared, and new model rollout trajectories and corresponding rewards are collected based on one epoch of data samples. For each data sample, the model generates $g$ trajectories, which are used to train the LoRA adapters over the subsequent 9 epochs. For each trajectory, we obtain its PDMS score vector $\textbf{r}=\{r_{1},r_{2},\dots,r_{g}\}$, enabling estimation of the advantage function (Shao et al., 2024):
@@ E.3 Inference and Evaluation
During inference, the DP-VLA encoder extracts features from the input images and driving commands. The denoising process is solved using the DPM-Solver(Lu et al., 2022), with the action decoder iteratively predicting the denoised trajectory over 10 steps to produce the final clean trajectory. For evaluation on the NAVSIM benchmark, the proposed trajectory is fed into an LQR tracker and dynamics model to compute the posterior trajectory. The final PDM score is derived from this posterior trajectory, satisfying the benchmark’s evaluation criteria (Dauner et al., 2024).